Using Maple to Bridge Between Calculus and “Freshman Engineering”

Douglas B. Meade

Department of Mathematics University of South Carolina

Columbia, SC 29208

(803) 777-6183

     meade@math.sc.edu
http://www.math.sc.edu/˜ meade/

20 November 1998

November 1998  ICTCM XI (New Orleans)  0

Using Maple to Bridge Between Calculus and “Freshman Engineering”

Douglas B. Meade

Department of Mathematics University of South Carolina

Columbia, SC 29208

(803) 777-6183

     meade@math.sc.edu
http://www.math.sc.edu/˜ meade/

20 November 1998

Outline

• Introduction

• Detailed Example: Water Quality • Other Examples

Introduction

• The Problem – as seen by Engineers

– as seen by Mathematicians

• A Solution – communicate and cooperate

November 1998  ICTCM XI (New Orleans)  3

Other Examples

• Strength and Toughness of Materials • Lift and Drag

• Bandwidth of an Optical Filter

• Low–Pass Electrical Filter

• Solvents and Solutes

— Engineer’s Toolkit: Maple V for Engineers,

D.B. Meade and E. Bourkoff, Addison Wesley Longman, 1998.